estion 8: Tree Trunks Co. makes two different types of pastries: Pies and Cakes. Tree Trunks Co. currently uses normal co
Answer #1
Pies Cakes
Direct Materials $         960,000 $         600,000
Direct Labor $         270,000 $         120,000
Manufacturing Overhead $         270,000 $         108,000
$     1,500,000 $         828,000
Units Produced             500,000             300,000
$                3.00 $                2.76

a. $ 3.00 per pie
b. $ 2.76 per cake

Here is the formula version

Pies Cakes Direct Materials Direct Labor Manufacturing Overhead 960000 270000 =178000/42000*30000 =SUM(D4:06) 500000 =D7/08 6

Activity Activity Measure Pies Cake Total Overhead Rate
General Factory Direct Labor Hours      30,000      12,000      42,000 $           84,000 $            2.00
Machine Use Machines Hours    200,000    100,000    300,000 $           90,000 $            0.30
Machine Setup Machine Set Ups            100            300            400 $         100,000 $       250.00
Distribution Units Produced    500,000    300,000    800,000 $         104,000 $            0.13
$         378,000
Pies Cakes
Direct Materials $         960,000 $         600,000
Direct Labor $         270,000 $         120,000
Manufacturing Overhead
General Factory $           60,000 $           24,000
Machine Use $           60,000 $           30,000
Machine Setup $           25,000 $           75,000
Distribution $           65,000 $           39,000
Total Cost $     1,440,000 $         888,000
Units Produced             500,000             300,000
$                2.88 $                2.96

c. $ 2.88
d. $ 2.96